julie blows a bubble. At first the pressure of the gas in the bubble is 4kPa. the bubble floats into the air and expands. when it gets to the top of a tree the bubble has a pressure of 2KPa and a volume of 3.5m^3. assuming a constant temperature, what was in the volume in m^3 of the bubble when it was first blown
1
Expert's answer
2020-06-16T09:29:17-0400

Since the temperature does not change, we can use the combined gas law to find the initial volume:


"\frac{P_1V_1}{T}=\frac{P_2V_2}{T},\\\space\\
V_1=V_2\frac{P_2}{P_1}=3.5\frac{2}{4}=1.75\text{ m}^3."